### v1

# Energetics-based model for a diffusiophoretic motion of a deformable droplet

Abstract
We construct a mathematical model for a diffusiophoretic motion of a deformable droplet, which is floating at a liquid surface and is driven by the surface tension gradient originating from the surface concentration field of the chemicals that are emitted from the droplet. We define the free energy of the system by including the surface and line energies. From the calculation of the functional of the free energy, we obtain a mathematical model for the diffusiophoretic motion with deformation. By only considering the deformation of the second mode, we explicitly derived the time-evolution equations for the translational motion and the elliptic deformation. We found three stable states: an immobile circular droplet, an immobile elliptically deformed droplet, and a moving droplet with the elliptic deformation in which the minor axis meets the motion direction, and discussed the transition between these three stable states.
preprint: APS/123-QED
I Introduction
Under nonequilibrium conditions, particles and droplets can spontaneously move through the dissipation of chemical energy, which is often called self-propulsion. They have been intensively studied as a model for the motion of living organisms. They have also attracted interest as elements of active matter, in which spatio-temporal order at greater scales than those of elements emerges [1, 2, 3]. Many experimental systems on self-propulsion in physico-chemical systems have also been reported; oil droplets in surfactant aqueous solutions or aqueous droplets in oil with surfactant exhibit self-propulsion [4, 5, 6, 7, 8, 9, 10, 11]. Other examples are Quincke rollers, self-propelled plastic particles under an electric field [12, 13], and Janus particles, two-faced particles that move due to the temperature gradient or chemical potential gradient [14, 15, 16, 17]. Floating objects on liquid surfaces can exhibit self-propulsion due to the difference in surface tension, which are often called Marangoni surfers [18, 19, 20, 21, 22]. Droplet motions with adsorption and desorption of surfactants were discussed based on the low-Reynolds number hydrodynamics [23, 24, 25], and the self-propelled droplet on the solid surface in a surfactant aqueous solution is discussed based on the energetics [26, 27, 26, 28]. The solid Marangoni surfers have also been discussed based on hydrodynamics [29, 30] or the coupled system between Newtonian equations and reaction-diffusion equations [31, 32, 19, 33].
In self-propulsion systems, the coupling between motion and shape has attracted interest. This is partly because the motion of living cells such as Keratocyte and Dictyostelium is strongly correlated with their shapes [34, 35, 36]. In also physico-chemical systems, the shape-dependent self-propulsion has been reported. Droplets that exhibit Marangoni surfing are often deformed and their shapes are strongly correlated with the direction of the motion [37, 38, 39, 40, 41, 42, 43, 44, 45]. This is because both self-propulsion and deformation have the same origin. For example, an oil droplet in a surfactant aqueous solution exhibits a self-propelled motion due to the Marangoni convection, which also causes ellipsoidal deformation [46]. A floating oil droplet at a water surface exhibits motion and deformation under a vertical vibration, both of which are led by Faraday waves [47]. Gel formation at the interface is also known to cause motion and deformation of the droplet [48] and a similar droplet behavior due to the actomyosin force generation was reported recently [49]. From a viewpoint of theoretical approach to the coupling between self-propulsion and deformation, there has also been a wide variety of mathematical modeling. One of the simplest models, so-called Ohta-Ohkuma model, described by coupled ordinary differential equations is proposed and analyzed based on the system symmetry [50, 51, 52]. This model only suggests the possible coupling manner and do not include the correspondence to the mechanism of the actual systems. On the other hand, some models have been developed based on insights from actual phenomena of the droplet motion and deformation [53, 54, 55]. Mathematical models that reproduce the motion and deformation of the cells have also been proposed [56, 57].
We have investigated the motion of Marangoni surfers based on experiments and mathematical modeling comprising the Newtonian equations and reaction-diffusion equations. The Marangoni surfers are driven by surface tension gradient originating from the concentration profile of chemicals that are emitted from the surfers. They tend to escape from the chemicals and thus their motion mechanism is often called self-diffusiophoresis. We have considered such a self-diffusiophoretic particle with a given shape that emits surface-active chemicals and moves due to the surface tension gradient [58, 59, 60, 61, 33]. For example, theoretical analysis suggests that an elliptic Marangoni surfer should move in its minor axis direction, which is also confirmed by numerical simulations and experiments using a camphor particle. In these studies, the surfer has a rigid shape, and thus we do not need to discuss the change in shape. In contrast, deformable droplets can also move as Marangoni surfers. For example, a pure alcohol droplet like pentanol and hexanol on an alcohol aqueous solution is known as the examples of such deformable moving droplets [37, 38, 62]. These experimental results clearly demonstrate the coupling between motion and deformation. Some of the authors proposed a mathematical model based on the energy minimization but the cost for numerical simulation was quite high [54]. Therefore, a mathematical model that well reproduces experimental results and enables us to discuss the mechanism from a viewpoint of system symmetry is awaited. Recently, some of the authors proposed a mathematical model adopting a phase-field description for the droplet shape [55]. The model succeeded in reproducing self-propulsion coupled with deformation but the relation to the simple models like the Ohta-Ohkuma model [50, 51], is still unveiled. Through the discussion of the correspondence between these models, we believe that we can attain the essential mechanism of the diffusiophoretic motion with deformation.
In the present study, thus, we propose a mathematical model for the diffusiophoretic deformable droplet, in which a droplet shape is described by the Fourier series expansion. Our model corresponds to a system with a floating deformable droplet that releases surface-active chemicals and generates their concentration field. The motion and deformation of the droplet are driven by the surface tension gradient originating from the concentration field of the chemicals. The system dynamics are partly derived using a variational principle of the free energy of the system. In the manuscript, we first show our mathematical model together with brief derivation of it in Sect. II. Then, we limit the case in which a droplet shows translational motion and the 2-mode deformation. We show the explicit form of the model and perform numerical simulations in Sect. III. Further, we perform theoretical analysis using the perturbation method and compare with the simulation results in Sect. IV. We discuss the results by numerical simulation and theoretical analyses in Sect. VI and summarize our work in Sect. VII. Since the present model is derived naturally from the free energy of the system, the model can be a universal one for the diffusiophoretic motion of a deformable object in a two-dimensional space.
II Modeling
We consider a two-dimensional system in which a deformable droplet is floating at a liquid surface and is driven by the surface tension gradient originating from the concentration profile of chemicals emitted from the droplet. It should be noted that we do not consider the hydrodynamics or depth of the system, but only consider a two-dimensional system. The deformation of a droplet is described by the deviation from a circular shape. We consider the local polar coordinates $(r,\theta)$ where the origin meets the center of mass $\bm{r}_{c}$ of the droplet. Then, the position of the droplet periphery is expressed as
$r=$ $f(\theta;\left\{a_{k}\right\},\left\{b_{k}\right\})$
$=$ $R\left[1+\sum_{k=2}^{\infty}\left(a_{k}\cos k\theta+b_{k}\sin k\theta\right)\right],$ (1)
where $R$ is the droplet radius. The coefficients $\left\{a_{k}\right\}$ and $\left\{b_{k}\right\}$ are on the order of a small parameter $\epsilon$, and indicate the deformation amplitudes for the $k$th mode. Here, we only consider the small deformation from a circle that can be described by the first order of the small parameter $\epsilon$. It is notable that the 0th and 1st modes are omitted since they correspond to the expansion/contraction and translation, respectively. The region inside the deformable droplet is given as
$\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})=\left\{\bm{r}\middle|\bm{r}-\bm{r}_{c}\in\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\})\right\},$ (2)
with
$\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\})=\left\{\bm{r}=r\bm{e}_{r}(\theta)\middle|r\leq f(\theta;\left\{a_{k}\right\},\left\{b_{k}\right\})\right\},$ (3)
where $\bm{e}_{r}(\theta)$ and $\bm{e}_{\theta}(\theta)$ are the unit vectors in the polar coordinates, which are explicitly defined as $\bm{e}_{r}(\theta)=\cos\theta\bm{e}_{x}+\sin\theta\bm{e}_{y}$ and $\bm{e}_{\theta}(\theta)=-\sin\theta\bm{e}_{x}+\cos\theta\bm{e}_{y}$ using the unit vectors in the Cartesian coordinates.
Considering that the droplet emits surface-active chemicals, and the emitted chemicals exhibit diffusion at the water surface and evaporate to the air phase or dissolve into a bulk aqueous phase, we consider the time evolution of the concentration $u$ of the chemicals obeys the reaction-diffusion equation as
$\frac{\partial u}{\partial t}=D\nabla^{2}u-\alpha u+S(\bm{r}-\bm{r}_{c};\left\{a_{k}\right\},\left\{b_{k}\right\}).$ (4)
Here, $D$ is an effective diffusion coefficient that includes the hydrodynamic effect [63, 64, 65], $\alpha$ is the evaporation and dissolution rate, and $S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})$ are the supply rate of the surface-active chemicals from the droplet. The term $S$ is explicitly expressed as
$S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})=\left\{\begin{array}[]{ll}S_{0}/A,&\bm{r}\in\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\}),\\ 0,&\bm{r}\notin\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\}).\end{array}\right$. (7)
Here, $S_{0}$ is the supply rate of the surface active chemicals from the droplet, and $A$ is the base area of the droplet, which is explicitly obtained as $A=\pi R^{2}+\mathcal{O}(\epsilon^{2})$.
In order to derive the time-evolution equations for the droplet velocity and deformation, we consider the free energy of the system as
$E=$ $E_{s}+E_{l}$ (8)
where the first term $E_{s}$ on the right-hand side denotes the surface energy depending on the surface tension $\gamma$ as a function of the local chemical concentration $u$ as
$E_{s}=\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma\left(u(\bm{r})\right)dA,$ (9)
and the second term $E_{l}$ denotes the line energy proportional to the peripheral length of the droplet,
$E_{l}=\tau\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}d\ell.$ (10)
Here, $dA$ and $d\ell$ denote the area and line elements, respectively, and $\partial\Omega$ denotes the periphery of the region $\Omega$. It is known that the line tension is negligibly small compared with the surface tension, but if the droplet is flat enough except for its periphery, it has extra energy which should be proportional to the peripheral length [66, 45]. We thus assume that $\tau$ is a positive constant.
The time evolution of the droplet velocity and deformation is given as
$\eta_{t}\frac{d\bm{r}_{c}}{dt}=-\frac{\partial E}{\partial\bm{r}_{c}},$ (11)
$\eta_{k}\frac{da_{k}}{dt}=-\frac{\partial E}{\partial a_{k}},$ (12)
$\eta_{k}\frac{db_{k}}{dt}=-\frac{\partial E}{\partial b_{k}},$ (13)
for $k=2,3,4,\dots$, where $\eta_{t}$ and $\eta_{k}$ are the positive coefficients to determine the time scale of the translational motion and deformation, respectively. The right sides of these equations can be calculated as
$\frac{\partial E_{s}}{\partial\bm{r}_{c}}$ $=-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla\gamma(u(\bm{r}))dA$
$=-\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{n}(\bm{r})d\ell,$ (14)
$\frac{\partial E_{s}}{\partial a_{k}}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell,$ (15)
$\frac{\partial E_{s}}{\partial b_{k}}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(b)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(b)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (16)
Here, we introduce incompressible vector fields $\bm{w}^{(a)}_{k}$ and $\bm{w}^{(b)}_{k}$, whose explicit expressions are given as
$\bm{w}^{(a)}_{k}$= $\nabla\left[\frac{R^{2}}{k}\left(\frac{r}{R}\right)^{k}\cos k\theta\right]$
$=$ $R\left(\frac{r}{R}\right)^{k-1}\left(\cos k\theta\bm{e}_{r}(\theta)-\sin k\theta\bm{e}_{\theta}(\theta)\right),$ (17)
$\bm{w}^{(b)}_{k}$= $\nabla\left[\frac{R^{2}}{k}\left(\frac{r}{R}\right)^{k}\sin k\theta\right]$
$=$ $R\left(\frac{r}{R}\right)^{k-1}\left(\sin k\theta\bm{e}_{r}(\theta)+\cos k\theta\bm{e}_{\theta}(\theta)\right).$ (18)
It should be noted that $\bm{w}_{k}^{(a)}$ and $\bm{w}_{k}^{(b)}$ correspond to the incompressible flow field that obeys the Stokes equation. The detailed calculation is shown in Appendix A.
As for the extra surface energy for the periphery, we obtain
$E_{l}$ $=2\pi\tau R\left[1+\frac{1}{4}\sum_{k=2}^{\infty}k^{2}\left({a_{k}}^{2}+{b_{k}}^{2}\right)\right]+\mathcal{O}(\epsilon^{3}).$ (19)
Therefore,
$\frac{\partial E_{l}}{\partial\bm{r}_{c}}=0,$ (20)
$\frac{\partial E_{l}}{\partial a_{k}}=\pi k^{2}\tau Ra_{k}+\mathcal{O}(\epsilon^{2})=\kappa_{k}a_{k}+\mathcal{O}(\epsilon^{2}),$ (21)
$\frac{\partial E_{l}}{\partial b_{k}}=\pi k^{2}\tau Rb_{k}+\mathcal{O}(\epsilon^{2})=\kappa_{k}b_{k}+\mathcal{O}(\epsilon^{2}),$ (22)
where $\kappa_{k}$ is defined as
$\kappa_{k}=\pi k^{2}\tau R>0.$ (23)
Thus, we finally obtain
$\eta_{t}\frac{d\bm{r}_{c}}{dt}$ $=\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla\gamma(u(\bm{r}))dA$
$=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{n}(\bm{r})d\ell,$ (24)
$\eta_{k}\frac{da_{k}}{dt}=$ $\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA-\kappa_{k}a_{k}$
$=$ $\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell-\kappa_{k}a_{k},$ (25)
$\eta_{k}\frac{db_{k}}{dt}=$ $\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(b)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA-\kappa_{k}b_{k}$
$=$ $\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(b)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell-\kappa_{k}b_{k}.$ (26)
The first line in each equation is suitable for the numerical simulation while the second line is suitable for the theoretical analysis.
The relationship between the surface tension and concentration of the chemical is assumed to be
$\gamma=\gamma_{0}-\Gamma u,$ (27)
where $\gamma_{0}$ is the surface tension of pure water and $\Gamma$ is a positive constant.
Hereafter, we consider the dimensionless version of the model, where the scales for length, time, concentration, and force are set as $\sqrt{D/\alpha}$, $1/\alpha$, $S_{0}/\alpha$, $\Gamma S_{0}/\sqrt{\alpha D}$, respectively. Then, the dimensionless version of our model is as follows
$\eta_{t}\frac{d\bm{r}_{c}}{dt}$ $=-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla u(\bm{r})dA$
$=-\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}u(\bm{r})\bm{e}_{n}(\bm{r})d\ell,$ (28)
$\eta_{k}\frac{da_{k}}{dt}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla u(\bm{r})dA-\kappa_{k}a_{k}$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}u(\bm{r})\bm{w}^{(a)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell-\kappa_{k}a_{k},$ (29)
$\eta_{k}\frac{db_{k}}{dt}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(b)}_{k}\cdot\nabla u(\bm{r})dA-\kappa_{k}b_{k},$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}u(\bm{r})\bm{w}^{(b)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell-\kappa_{k}b_{k},$ (30)
$\frac{\partial u}{\partial t}=\nabla^{2}u-u+S(\bm{r}-\bm{r}_{c};\left\{a_{k}\right\},\left\{b_{k}\right\}),$ (31)
$S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})=\left\{\begin{array}[]{ll}1/A,&\bm{r}\in\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\}),\\ 0,&\bm{r}\notin\Omega_{0}(\left\{a_{k}\right\},\left\{b_{k}\right\}),\end{array}\right$. (34)
with Eqs. (2) and (3).
III Numerical simulation in the case with 2-mode deformation
Based on the model derived above, we performed a numerical simulation. For the correspondence to the theoretical analyses described in the following section, we only consider the 2-mode deformation. In order to minimize the effect of discretization, we define a smooth function representing the area inside the droplet as
$\Theta_{\sigma}(\bm{r};a_{2},b_{2})=$ $\frac{1}{2}\left[1+\tanh\left(-\frac{r-f(\theta;a_{2},b_{2})}{\sigma}\right)\right],$ (35)
where
$f(\theta;a_{2},b_{2})=R\left(1+a_{2}\cos 2\theta+b_{2}\sin 2\theta\right).$ (36)
Here, $\sigma$ is a positive constant for the smooth connection of variables at the boundary. As for the calculation of the time evolution of $\bm{r}_{c}$, $a_{2}$, and $b_{2}$, we adopt $\Theta_{\sigma}(\bm{r}_{c};a_{2},b_{2})$ as
$\eta_{t}\frac{d\bm{r}_{c}}{dt}=-\iint_{\mathbb{R}^{2}}\nabla u(\bm{r})\Theta_{\sigma}(\bm{r}-\bm{r}_{c};a_{2},b_{2})dA,$ (37)
$\eta_{2}\frac{da_{2}}{dt}=-\iint_{\mathbb{R}^{2}}\bm{w}^{(a)}_{2}\cdot\nabla u(\bm{r})\Theta_{\sigma}(\bm{r}-\bm{r}_{c};a_{2},b_{2})dA-\kappa_{2}a_{2},$ (38)
$\eta_{2}\frac{db_{2}}{dt}=$ $-\iint_{\mathbb{R}^{2}}\bm{w}^{(b)}_{2}\cdot\nabla u(\bm{r})\Theta_{\sigma}(\bm{r}-\bm{r}_{c};a_{2},b_{2})dA-\kappa_{2}b_{2},$ (39)
$\frac{\partial u}{\partial t}=\nabla^{2}u-u+S_{\sigma}(\bm{r}-\bm{r}_{c};a_{2},b_{2}),$ (40)
where
$S_{\sigma}(\bm{r};a_{2},b_{2})=\frac{1}{A}\Theta_{\sigma}(\bm{r};a_{2},b_{2}).$ (41)
It should be noted that $S_{\sigma}(\bm{r};a_{2},b_{2})$ converges to $S(\bm{r};\left\{a_{k}\right\},\left\{b_{k}\right\})$ in Eq. (34) with $a_{3}=a_{4}=\dots=0$ and $b_{3}=b_{4}=\dots=0$ when $\sigma\to+0$. $\bm{w}_{2}^{(a)}$ and $\bm{w}_{2}^{(b)}$ are explicitly given as
$\bm{w}_{2}^{(a)}=x\bm{e}_{x}-y\bm{e}_{y},$ (42)
$\bm{w}_{2}^{(b)}=y\bm{e}_{x}+x\bm{e}_{y}.$ (43)
For the numerical simulation, we adopted an explicit method for the time development with spatial mesh $\Delta x=0.025$ and time step $\Delta t=0.0001$. The system size was set to be $7\times 7$ with the periodic boundary condition. The smoothing parameter was set to be $\sigma=0.03$. We checked that the results did not change significantly if we slightly changed the value of $\sigma$. We fixed $R=1$ and $\eta_{2}=0.1$, and varied $\eta_{t}$ and $\kappa_{2}$ as the parameters. The initial conditions for $u$ and $\bm{r}_{c}$ were fixed as $u=0$ and $\bm{r}_{c}=\bm{0}$, and those for $d\bm{r}_{c}/dt$, $a_{2}$, and $b_{2}$ were varied.
Figure 1: Snapshots at $t=1000$ (left panels) and time series (right panels) of the speed $v=\left|d\bm{r}_{c}/dt\right|$ (red) and the deformation magnitude $s=\sqrt{{a_{2}}^{2}+{b_{2}}^{2}}$ (blue) for the three typical cases. (a) Immobile circular (IC) droplet at $\kappa_{2}=0.12$ and $\eta_{t}=0.085$. (b) Immobile deformed (ID) droplet at $\kappa_{2}=0.116$ and $\eta_{t}=0.085$. (c) Mobile deformed (MD) droplet at $\kappa_{2}=0.112$ and $\eta_{t}=0.085$. The droplet moves downward as the orange arrow indicates. Grayscale tone displays the concentration field $u$. The initial condition was set as $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$.
In Fig. 1, we show a typical example of the droplet dynamics. For great $\eta_{t}$ and $\kappa_{2}$, we observed an immobile circular (IC) droplet, which kept a circular shape and did not move as shown in Fig. 1(a). For smaller $\kappa_{2}$ and greater $\eta_{t}$, we observed an immobile deformed (ID) droplet, which deformed into an elliptic or peanut shape but did not move, as shown in Fig. 1(b). For smaller $\eta_{t}$, we observed a moving deformed (MD) droplet, which moved in a certain direction with an elliptic deformation whose direction of the minor axis met the direction of motion as shown in Fig. 1(c).
Figure 2: Phase diagram on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters were set to be $\eta_{2}=0.1$ and $R=1$. The red, green, and blue points correspond to an immobile circular (IC) droplet, an immobile deformed (ID) droplet, and a mobile deformed (MD) droplet. The cyan points show the bistability between IC and MD droplets. The orange points show the other cases, in which we could classify into none of the above.
In Fig. 2, we show a phase diagram of stable states on the $\eta_{t}$-$\kappa_{2}$ plane. The other parameters were fixed at $\eta_{2}=0.1$ and $R=1$. We calculated with the following four initial conditions (i) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.1\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, (ii) $d\bm{r}_{c}/dt=\bm{e}_{x}+0.01\bm{e}_{y}$, $a_{2}=-0.25$, $b_{2}=0.002$, (iii) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.1$, $b_{2}=0.01$, and (iv) $d\bm{r}_{c}/dt=0.01\bm{e}_{x}+0.001\bm{e}_{y}$, $a_{2}=0.01$, $b_{2}=0.001$, where the initial conditions (i) and (ii) have a large velocity and deformation with different direction relations, (iii) has a small velocity and a large deformation, and (iv) has a small velocity and deformation. In the phase diagram, the IC droplet was stable for $\eta_{t}\gtrsim 0.077$ and $\kappa_{2}\gtrsim 0.119$. With a decrease in $\kappa_{2}$, the IC droplet became unstable and an ID droplet became stable. In contrast, with a decrease in $\eta_{t}$, an IC droplet became unstable and the MD droplet was realized. Close to the transition points between IC and MD droplets, we observed the bistable state of them. The orange points around $\eta_{t}\simeq 0.077$ and $\kappa_{2}\gtrsim 0.153$ had a finite speed with a small deformation. They could be classified into another type like “mobile circular droplets”, but since we only observed them at few parameter values and found they showed slight deformation, we showed them as “cannot be classified”.
In Fig. 3, the one-dimensional bifurcation diagrams are shown. Here, we adopted symmetric initial conditions (i’) $d\bm{r}_{c}/dt=\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0.0$, and (ii’) $d\bm{r}_{c}/dt=0.001\bm{e}_{x}$, $a_{2}=-0.1$, $b_{2}=0$ to clearly show the bifurcation structures of the solutions. In Fig. 3(a), the bifurcation diagram with respect to $\kappa_{2}$ at $\eta_{t}=0.085$ is shown. The deformation magnitude $s$ rose at $\kappa_{2}\lesssim 0.118$. For $0.115\lesssim\kappa_{2}\lesssim 0.118$, the droplet speed was zero, which indicated an ID droplet. For $\kappa_{2}\lesssim 0.115$, another branch with finite speed and deformation amplitude discontinuously appeared, which indicated an MD droplet. The transition from an IC droplet to an ID droplet at $\kappa_{2}\simeq 0.118$ seems to be a supercritical pitchfork bifurcation. In contrast, the transition from an ID droplet to an MD droplet seems discontinuous since there exists a bistable state. In Fig. 3(b), the bifurcation diagram with respect to $\eta_{t}$ for $\kappa_{2}=0.13$ is shown. The droplet speed $v$ and the deformation amplitude simultaneously rose at $\eta_{t}\lesssim 0.079$. The IC droplet became unstable at $\eta_{t}\simeq 0.076$. The transition from an IC droplet to an MD droplet seemed discontinuous, which implies a subcritical pitchfork bifurcation.
Figure 3: Phase diagrams with constant $\eta_{t}$ and $\kappa_{2}$, which correspond to the vertical and horizontal lines, respectively. (a) Phase diagram at $\eta_{t}=0.085$. Speed $v$ and deformation $s$ are plotted depending on $\kappa_{2}$. (b) Phase diagram at $\kappa_{2}=0.13$ depending on $\eta_{t}$. Speed $v$ and deformation $s$ are plotted depending on $\kappa_{2}$. The red, blue, and green points represent immobile circular (IC), immobile deformed (ID), and mobile deformed (MD) droplets. The vertical dotted lines are drawn to indicate the transition points and correspondence between $v$ and $s$.
For even smaller $\eta_{t}$ or $\kappa_{2}$, the calculation diverged due to $s$. This is because our modeling is based on the assumption that the deformation is sufficiently small as the first-order perturbation from a circle. Actually, $s$ is well-defined in the range of $\left|s\right|<1$.
IV Reduction to ODE model including only 2-mode deformation
Here, we also consider a droplet motion only with the 2-mode deformation. The driving force for translational motion $\bm{F}$ and those for deformation $G^{(2a)}$ and $G^{(2b)}$ that originate from $E_{s}$ are defined as
$\bm{F}=-\oint_{\partial\Omega\left(\bm{r}_{c},a_{2},b_{2}\right)}u(\bm{r})\bm{e}_{n}(\bm{r})d\ell,$ (44)
$G^{(2a)}=-\oint_{\partial\Omega\left(\bm{r}_{c},a_{2},b_{2}\right)}u(\bm{r})\bm{w}^{(a)}_{2}\cdot\bm{e}_{n}(\bm{r})d\ell,$ (45)
$G^{(2b)}=-\oint_{\partial\Omega\left(\bm{r}_{c},a_{2},b_{2}\right)}u(\bm{r})\bm{w}^{(b)}_{2}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (46)
We first calculate the steady-state concentration field in the case that the droplet is moving at a constant velocity $v\bm{e}_{x}$ and constant deformation amplitudes $a_{2}$ and $b_{2}$ as shown in Appendix B. Then, we plug it into Eqs. (44), (45), and (46) to obtain the explicit forms for $\bm{F}$, $G^{(2a)}$, and $G^{(2b)}$ as
$\bm{F}=$ $\left[\left(F^{(x)}_{1}+\tilde{F}^{(x)}_{1}a_{2}\right)v+\left(F^{(x)}_{3}+\tilde{F}^{(x)}_{3}a_{2}\right)v^{3}\right]\bm{e}_{x}$
$+\left[\tilde{F}^{(y)}_{1}b_{2}v+\tilde{F}^{(y)}_{3}b_{2}v^{3}\right]\bm{e}_{y},$ (47)
$G^{(2a)}=G_{2}^{(2a)}v^{2}+\tilde{G}^{(2a)}_{0}a_{2}+\tilde{G}_{2}^{(2a)}a_{2}v^{2},$ (48)
$G^{(2b)}=\tilde{G}^{(2b)}_{0}b_{2}+\tilde{G}^{(2b)}_{2}b_{2}v^{2},$ (49)
where
$F^{(x)}_{1}=\frac{\pi R^{4}}{4A}\left[\mathcal{I}_{0}\left(R\right)\mathcal{K}_{0}\left(R\right)-\mathcal{I}_{2}\left(R\right)\mathcal{K}_{2}\left(R\right)\right],$ (50)
$F^{(x)}_{3}$= $\frac{\pi R^{6}}{32A}\left[\mathcal{I}_{0}(R)\mathcal{K}_{0}(R)-\frac{2}{R^{2}}\mathcal{I}_{1}(R)\mathcal{K}_{1}(R)\right$.
$\left.\qquad-\mathcal{I}_{2}(R)\mathcal{K}_{2}(R)\right],$ (51)
$\tilde{F}^{(x)}_{1}=\tilde{F}^{(y)}_{1}=-\frac{\pi R^{4}}{2A}\left[\mathcal{I}_{1}\left(R\right)\mathcal{K}_{1}\left(R\right)-\mathcal{I}_{2}\left(R\right)\mathcal{K}_{2}\left(R\right)\right],$ (52)
$\tilde{F}^{(x)}_{3}=2\tilde{F}^{(y)}_{3}$= $\frac{\pi R^{6}}{48A}\left[3\mathcal{I}_{0}\left(R\right)\mathcal{K}_{0}\left(R\right)-4\mathcal{I}_{1}\left(R\right)\mathcal{K}_{1}\left(R\right)\right$.
$\left.\qquad+\mathcal{I}_{2}\left(R\right)\mathcal{K}_{2}\left(R\right)\right],$ (53)
$G^{(2a)}_{2}$= $\frac{\pi R^{6}}{64A}\left[-2\mathcal{I}_{0}(R)\mathcal{K}_{0}(R)+\mathcal{I}_{1}(R)\mathcal{K}_{1}(R)\right$.
$\left.\qquad+2\mathcal{I}_{2}(R)\mathcal{K}_{2}(R)-\mathcal{I}_{3}(R)\mathcal{K}_{3}(R)\right],$ (54)
$\tilde{G}^{(2a)}_{0}=\tilde{G}^{(2b)}_{0}=\frac{\pi R^{4}}{A}\left[\mathcal{I}_{1}\left(R\right)\mathcal{K}_{1}\left(R\right)-\mathcal{I}_{2}\left(R\right)\mathcal{K}_{2}\left(R\right)\right],$ (55)
$\tilde{G}^{(2a)}_{2}=\tilde{G}^{(2b)}_{2}$= $\frac{\pi R^{6}}{32A}\left[-4\mathcal{I}_{0}(R)\mathcal{K}_{0}(R)+7\mathcal{I}_{1}(R)\mathcal{K}_{1}(R)\right$.
$\left.\qquad-4\mathcal{I}_{2}(R)\mathcal{K}_{2}(R)+\mathcal{I}_{3}(R)\mathcal{K}_{3}(R)\right].$ (56)
It is notable that $F^{(x)}_{1}>0$, $F^{(x)}_{3}<0$, $\tilde{F}^{(x)}_{1}=\tilde{F}^{(y)}_{1}<0$, $G_{2}^{(2a)}<0$, $\tilde{G}_{0}^{(2a)}=\tilde{G}_{0}^{(2b)}>0$ for all $R>0$. The other coefficients $\tilde{F}_{3}^{(x)}$, $\tilde{F}_{3}^{(y)}$, $\tilde{G}_{2}^{(2a)}$, and $\tilde{G}_{2}^{(2b)}$ change their signs depending on $R$.
In order to perform analyses based on dynamical systems, we introduce an inertia term to the equation of translational motion in Eq. (28) as
$m\frac{d\bm{v}_{c}}{dt}=-\eta_{t}\bm{v}_{c}+\bm{F}.$ (57)
The addition of the inertia term $m\,d\bm{v}_{c}/dt$ with effective mass $m$ is justified by the following two reasons. The first is that the particle should have a mass even if it is small, and the second is that the inertia term naturally appears by the expansion of the concentration field with respect to the time derivative of the velocity [67, 68].
V Analysis of ODE model
In the previous section, the analysis is made under the constraint that the velocity is in the positive $x$-axis direction. Taking rotational symmetry of the system into consideration, the evolution equation is given as
$\eta_{t}\frac{d\bm{v}_{c}}{dt}=$ $(-\eta_{t}+f_{1})\bm{v}_{c}-\tilde{f}_{1}S\bm{v}_{c}-f_{3}\left|\bm{v}_{c}\right|^{2}\bm{v}_{c}$
$+\tilde{f}_{3}\left\{\left|\bm{v}_{c}\right|^{2}S\bm{v}_{c}+\left(\bm{v}_{c}\cdot S\bm{v}_{c}\right)\bm{v}_{c}\right\},$ (58)
$\eta_{2}\frac{dS}{dt}=\left(-\kappa_{2}+\tilde{g}_{0}\right)S-g_{2}\left(2\bm{v}_{c}\otimes\bm{v}_{c}-\left|\bm{v}_{c}\right|^{2}I\right)-\tilde{g}_{2}\left|\bm{v}_{c}\right|^{2}S.$ (59)
Here, $I$ is an identity tensor and $S$ is a traceless tensor that represents the second-order deformation, which is connected with $a_{2}$ and $b_{2}$ [50, 52] as
$S_{xx}=-S_{yy}=a_{2},$ (60)
$S_{xy}=S_{yx}=b_{2}.$ (61)
The coefficients in Eqs. (58) and (59) are given by Eqs. (50) to (56) as $f_{1}=F^{(x)}_{1}$, $\tilde{f}_{1}=-\tilde{F}^{(x)}_{1}=-\tilde{F}^{(y)}_{1}$, $f_{3}=-F^{(x)}_{3}$, $\tilde{f}_{3}=\tilde{F}_{3}^{(x)}/2=\tilde{F}_{3}^{(y)}$, $\tilde{g}_{0}=\tilde{G}_{0}^{(2a)}=\tilde{G}_{0}^{(2b)}$, $g_{2}=-G^{(2a)}_{2}$, and $\tilde{g}_{2}=-\tilde{G}_{2}^{(2a)}=-\tilde{G}_{2}^{(2b)}$. These coefficients have positive values ($f_{1}=0.07812$, $\tilde{f}_{1}=0.05980$, $f_{3}\simeq 0.01150$, $\tilde{f}_{3}\simeq 0.09292$, $\tilde{g}_{0}\simeq 0.1196$, $g_{2}\simeq 0.006909$, $\tilde{g}_{2}\simeq 0.01487$) when $R=1$, which corresponds to the numerical simulation in Section III.
Considering the terms up to the third order of small parameters $\bm{v}_{c}$ and $S$, we obtain
$m\frac{d\bm{v}_{c}}{dt}=\left(-\eta_{t}+f_{1}\right)\bm{v}_{c}-\tilde{f}_{1}S\bm{v}_{c}-f_{3}\left|\bm{v}_{c}\right|^{2}\bm{v}_{c},$ (62)
$\eta_{2}\frac{dS}{dt}=$ $\left(-\kappa_{2}+\tilde{g}_{0}\right)S-g_{2}\left(2\bm{v}_{c}\otimes\bm{v}_{c}-\left|\bm{v}_{c}\right|^{2}I\right)$
$-\tilde{g}_{2}\left|\bm{v}_{c}\right|^{2}S-h_{0}S^{3}.$ (63)
It should be noted that here we phenomenologically add the term $h_{0}S^{3}$ in order to realize stable deformation of the droplet, though the term proportional to $S^{3}$ does not appear from the expansion in the discussion above since we only consider the first order of the deformation amplitude. Actually, the value of $h_{0}$ in the case of $R=1$, corresponding to the numerical simulation in Section III, can be determined by the fitting and it is given as $h_{0}\simeq 0.1436$ (see the details in Appendix C).
For the stability analysis, we consider in the polar coordinates as
$\bm{v}_{c}=v\left(\cos\phi\,\bm{e}_{x}+\sin\phi\,\bm{e}_{y}\right),$ (64)
$S=s\begin{pmatrix}\cos 2\psi&\sin 2\psi\\ \sin 2\psi&-\cos 2\psi\end{pmatrix}=\begin{pmatrix}a_{2}&b_{2}\\ b_{2}&-a_{2}\end{pmatrix}$. (65)
It should be noted that the angle $\psi$ indicates the elongated direction of the 2-mode deformation from a circle and the deformation amplitude is denoted by $s$. Here, we calculate $S^{3}$ as
$(S^{3})_{\alpha\beta}=S_{\alpha\gamma}S_{\gamma\delta}S_{\delta\beta}=s^{2}\delta_{\alpha\delta}S_{\delta\beta}=s^{2}S_{\alpha\beta}.$ (66)
Therefore, we finally obtain the ordinary differential equation (ODE) model as
$m\frac{dv}{dt}=\left(-\eta_{t}+f_{1}\right)v-\tilde{f}_{1}sv\cos 2(\psi-\phi)-f_{3}v^{3},$ (67)
$m\frac{d\phi}{dt}=-\tilde{f}_{1}s\sin 2(\psi-\phi),$ (68)
$\eta_{2}\frac{ds}{dt}=\left(-\kappa_{2}+\tilde{g}_{0}\right)s-g_{2}v^{2}\cos 2(\psi-\phi)-\tilde{g}_{2}v^{2}s-h_{0}s^{3},$ (69)
$\eta_{2}\frac{d\psi}{dt}=\frac{1}{2}g_{2}\frac{v^{2}}{s}\sin 2(\psi-\phi).$ (70)
Based on the ODE model, we performed the analyses of the solutions and their stability. The details are shown in Appendix D.
We adopted the coefficient values that corresponded to the numerical simulation ($R=1$) in Section III and calculated the phase diagram shown in Fig. 4. The diagram well reproduces the one obtained from the direct numerical simulation shown in Fig. 2. The slight differences should be due to the effect of the smoothing of the surface and space discretization. The most serious difference is that the bistable region between ID and MD droplets exists in the phase diagram in Fig. 4, while it does not exist in the one in Fig. 2. We consider this is because we neglect the higher-order terms in the process of reduction to the ODE model.
Figure 4: Phase diagram obtained by the ODE model. The boundary curves are given as $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (112) with Eq. (111), and Eq. (114). IC, ID, and MD indicate an immobile circular droplet, an immobile deformed droplet, and a mobile deformed droplet, respectively.
Therefore, our theoretical analyses on the bifurcation structure and numerical simulation results suggest that an IC droplet becomes unstable and a minor-axis directed moving elliptic droplet becomes stable with decreasing $\eta_{t}$ through a subcritical pitchfork bifurcation. In contrast, an IC droplet becomes unstable and an ID droplet becomes stable with decreasing $\kappa_{2}$ through a supercritical pitchfork bifurcation. With a further decrease in $\kappa_{2}$, an ID droplet becomes unstable and an MD droplet becomes stable. This transition is discontinuous but we could not identify the bifurcation structure since our reduced ODE model does not properly reflect the transition observed in direct numerical simulation.
VI Discussion
In the present paper, we constructed a model for a deformable self-propelled droplet that exhibits self-diffusiophoresis and performed numerical simulations and theoretical analyses to discuss the relation between motion and deformation using the Fourier expansion for the deformation. The analyses based on the Fourier mode expansion were also reported in previous studies. Ohta and Ohkuma proposed a mathematical model in which the velocity of the self-propelled particle and the amplitude of 2-mode deformation are considered [50, 52]. In their model, it is assumed that deformation decays if the particle does not move. In contrast, we derived the coefficients appearing in the Ohta-Ohkuma model by the reduction from a model including a concentration profile. The obtained results suggest both spontaneous translational motion and spontaneous deformation are possible. Especially, they did not discuss an immobile deformed (ID) droplet since they only considered the case with the decay of deformation. Here, we propose that the ID droplet can exist.
As for the pulse dynamics in reaction-diffusion systems, there have been several studies in which the coupling between the pulse shape and motion is concerned. Krischer and Mikhailov reported a traveling pulse in a two-dimensional reaction-diffusion system with volume conservation using numerical simulation. They exhibited a characteristic shape of a traveling solution [69]. Teramoto et al. studied the reaction-diffusion system with one activator and two inhibitors. They analyzed the bifurcation structure based on the center manifold theorem. They clarified the bifurcation structure including the bifurcation from a circular-shaped standing pulse into a deformed traveling pulse solution [70]. Ohta et al. analyzed the reaction-diffusion equation with volume conservation and discussed the dynamics by expanding with Fourier series [71]. They also expanded their study into a three-dimensional system [72]. Our approach is similar to their work in that the shape is described as the Fourier expansion in polar coordinates. Since our reaction-diffusion equation is linear, the analysis is simpler and the analytical approach can be more easily performed.
Recently, some of the authors published a paper on the phase-field model of a self-propelled droplet almost in the same mechanism, in which a droplet shape is described by a phase field and the evolution equation was derived by the variational principle of free energy. There, the relation between the time evolution of the phase field and that of the interface was discussed [55]. This work can be considered to be the case where the droplet shape is described as a function of the angle and the function is expanded as the Fourier series. Thanks to this simplification, we can discuss the stability of motion in our present model, though we could only perform numerical simulations to obtain the droplet motion in our previous model. We hope the universal understanding of a self-propelled deformed droplet will be obtained by the comparison between the previous phase-field model and the present model using the Fourier expansion.
VII Conclusion
We constructed a mathematical model for a deformable diffusio-phoretic self-propelled droplet, which moves by the surface tension gradient originating from the concentration field of the chemicals emitted from the droplet. By defining the free energy of the system including the surface and line energies, and calculating the variation of it, we obtained the time-evolution equations for the translational motion and deformation. In the case only including the 2-mode deformation, we performed numerical simulations and theoretical analyses, and obtained that an immobile deformed droplet or a mobile deformed droplet that moves in its minor-axis direction becomes stable through a supercritical and subcritical pitchfork bifurcation from an immobile circular droplet, respectively.
Acknowledgements.
This work was supported by JSPS KAKENHI Grants Nos. JP22K03428, JP23K20815, JP23H04936, JP24K06978, JP24K22311, JP24K16981, JP25K00918, and the Cooperative Research Program of “Network Joint Research Center for Materials and Devices” (Nos. 20254003, 20251011). This work was also supported by JSPS and PAN under the Japan-Poland Research Cooperative Program (No. JPJSBP120204602). This work was supported by the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
Appendix A Detailed derivation of $\partial E/\partial\bm{r}_{c}$, $\partial E/\partial a_{k}$, and $\partial E/\partial b_{k}$ in Eqs. (14), (15), and (16)
Here, we show the detailed derivation of Eqs. (14) , (15), and (16). We set $\bm{r}_{c}=x_{c}\bm{e}_{x}+y_{c}\bm{e}_{y}$, and first we calculate $\partial E_{s}/\partial x_{c}$ as
$\frac{\partial E_{s}}{\partial x_{c}}=$ $\lim_{\delta\to 0}\frac{\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c}+\delta\bm{e}_{x},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA-\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $\lim_{\delta\to 0}\frac{\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}+\delta\bm{e}_{x}))dA-\iint_{\mathbb{R}^{2}\backslash\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $-\lim_{\delta\to 0}\frac{\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}+\delta\bm{e}_{x}))dA-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $-\lim_{\delta\to 0}\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\frac{\gamma(u(\bm{r}+\delta\bm{e}_{x}))-\gamma(u(\bm{r}))}{\delta}dA$
$=$ $-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\frac{\partial\gamma(u(\bm{r}))}{\partial x}dA.$ (71)
In the same manner, we obtain
$\frac{\partial E_{s}}{\partial y_{c}}=$ $-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\frac{\partial\gamma(u(\bm{r}))}{\partial y}dA.$ (72)
and therefore we get
$\frac{\partial E_{s}}{\partial\bm{r}_{c}}=$ $-\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla\gamma(u(\bm{r}))dA.$ (73)
Using the Gauss’ divergence law,
$\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\frac{\partial\gamma(u(\bm{r}))}{\partial x}dA$ $=\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\bm{e}_{x}\cdot\nabla\gamma(u(\bm{r}))dA$
$=\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla\cdot(\gamma(u(\bm{r}))\bm{e}_{x})dA$
$=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{x}\cdot\bm{e}_{n}(\bm{r})d\ell,$ (74)
and in the same manner
$\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\frac{\partial\gamma(u(\bm{r}))}{\partial y}dA$ $=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{y}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (75)
Therefore,
$\iint_{\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\nabla\gamma(u(\bm{r}))dA$ $=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\left[(\bm{e}_{x}\cdot\bm{e}_{n}(\bm{r}))\bm{e}_{x}+(\bm{e}_{y}\cdot\bm{e}_{n}(\bm{r}))\bm{e}_{y}\right]d\ell$
$=\oint_{\partial\Omega(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\})}\gamma(u(\bm{r}))\bm{e}_{n}(\bm{r})d\ell,$ (76)
and we obtain Eq. (14).
Next, we calculate $\partial E_{s}/\partial a_{k}$. We define $\left\{e_{n}^{(k)}\right\}$ as
$e_{n}^{(k)}=\delta_{nk},$ (77)
where $\delta_{ij}$ is the Kronecker’s delta. Using this, we obtain
$\frac{\partial E_{s}}{\partial a_{k}}=$ $\lim_{\delta\to 0}\frac{\iint_{\mathbb{R}^{2}\backslash\Omega\left(\bm{r}_{c},\left\{a_{k}\right\}+\delta\left\{e_{n}^{(k)}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))dA-\iint_{\mathbb{R}^{2}\backslash\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $-\lim_{\delta\to 0}\frac{\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\}+\delta\left\{e_{n}^{(k)}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))dA-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $-\lim_{\delta\to 0}\frac{\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}+\delta\bm{w}^{(a)}_{k}))dA-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))dA}{\delta}$
$=$ $-\lim_{\delta\to 0}\frac{1}{\delta}\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\left[\gamma(u(\bm{r}+\delta\bm{w}^{(a)}_{k}))-\gamma(u(\bm{r}))\right]dA$
$=$ $-\lim_{\delta\to 0}\frac{1}{\delta}\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\left[\delta\bm{w}^{(a)}_{k}\cdot\nabla\gamma(u(\bm{r}))+\mathcal{O}(\delta^{2})\right]dA$
$=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA.$ (78)
In the calculation from the second to third lines, we replace $\bm{r}$ with $\bm{r}+\delta\bm{w}_{k}^{(a)}$, where $\bm{w}_{k}^{(a)}$ is given in Eq. (17). The Jacobian matrix regarding this transform is $1+\mathcal{O}(\epsilon^{2})$. The conditions required for the vector field $\bm{w}_{k}^{(a)}$ are that the divergence is 0, i.e., $\nabla\cdot\bm{w}_{k}^{(a)}=0$ and that it satisfies
$f(\theta,\{a_{k}\}+\delta\{e_{n}^{(k)}\},\{b_{k}\})\bm{e}_{r}=f(\theta,\{a_{k}\},\{b_{k}\})\bm{e}_{r}+\delta\bm{w}^{(a)}_{k}|_{\bm{r}=f(\theta,\{a_{k}\},\{b_{k}\})\bm{e}_{r}}.$ (79)
The incompressible flow profile in Eq. (17) holds the above conditions.
Using the Gauss’ divergence theorem, we obtain
$\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla\gamma(\bm{r})dA=$ $\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\nabla\cdot\left(\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\right)dA$
$=$ $\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (80)
Therefore, we obtain
$\frac{\partial E_{s}}{\partial a_{k}}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(a)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(a)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell.$ (81)
In the same manner, we obtain
$\frac{\partial E_{s}}{\partial b_{k}}=$ $-\iint_{\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\bm{w}^{(b)}_{k}\cdot\nabla\gamma(u(\bm{r}))dA$
$=$ $-\oint_{\partial\Omega\left(\bm{r}_{c},\left\{a_{k}\right\},\left\{b_{k}\right\}\right)}\gamma(u(\bm{r}))\bm{w}^{(b)}_{k}\cdot\bm{e}_{n}(\bm{r})d\ell,$ (82)
and thus we get Eqs. (15) and (16).
Appendix B Detailed derivation of the steady-state concentration
Here, we show the steady-state concentration $u_{s}$ when the droplet is moving at a constant velocity $\bm{v}_{c}$ and constant deformation amplitudes $a_{2}$ and $b_{2}$. For simplicity, we assume $\bm{v}_{c}=v\bm{e}_{x}$. The concentration field should hold
$-v\frac{\partial u_{s}}{\partial x}=\frac{\partial^{2}u_{s}}{\partial x^{2}}+\frac{\partial^{2}u_{s}}{\partial y^{2}}-u_{s}+S(\bm{r};\bm{r}_{c},a_{2},b_{2}),$ (83)
with
$S(\bm{r};\bm{r}_{c},a_{2},b_{2})=\left\{\begin{array}[]{ll}1/A,&\bm{r}\in\Omega(\bm{r}_{c},a_{2},b_{2}),\\ 0,&\bm{r}\notin\Omega(\bm{r}_{c},a_{2},b_{2}).\end{array}\right$. (86)
and
$\Omega(\bm{r}_{c},a_{2},b_{2})=\left\{\bm{r}=\bm{r}_{c}+r\bm{e}_{r}(\theta)\middle|r\leq R\left(1+a_{2}\cos 2\theta+b_{2}\sin 2\theta\right)\right\}.$ (87)
We also need to consider the continuity condition at the droplet boundary. The results are already given in our previous papers [58, 59, 60] on a self-propelled solid particle with a deformed shape from a circle. Here, we set $a_{2}=\epsilon\tilde{a}_{2}$ and $b_{2}=\epsilon\tilde{b}_{2}$ to clearly show the dependence of the small parameter $\epsilon$. Then, we obtain
$u_{s}=\left\{\begin{array}[]{ll}u_{s}^{(\mathrm{i})}=u^{(\mathrm{i})}_{0}+u^{(\mathrm{i})}_{1}v+u^{(\mathrm{i})}_{2}v^{2}+u^{(\mathrm{i})}_{3}v^{3}+\epsilon\left(\tilde{u}^{(\mathrm{i})}_{0}+\tilde{u}^{(\mathrm{i})}_{1}v+\tilde{u}^{(\mathrm{i})}_{2}v^{2}+\tilde{u}^{(\mathrm{i})}_{3}v^{3}\right)+\mathcal{O}(v^{4},\epsilon^{2}),&\bm{r}\in\Omega(\bm{r}_{c},a_{2},b_{2}),\\ u_{s}^{(\mathrm{o})}=u^{(\mathrm{o})}_{0}+u^{(\mathrm{o})}_{1}v+u^{(\mathrm{o})}_{2}v^{2}+u^{(\mathrm{o})}_{3}v^{3}+\epsilon\left(\tilde{u}^{(\mathrm{o})}_{0}+\tilde{u}^{(\mathrm{o})}_{1}v+\tilde{u}^{(\mathrm{o})}_{2}v^{2}+\tilde{u}^{(\mathrm{o})}_{3}v^{3}\right)+\mathcal{O}(v^{4},\epsilon^{2}),&\bm{r}\notin\Omega(\bm{r}_{c},a_{2},b_{2}),\end{array}\right$. (90)
where the explicit forms are given in the polar coordinates as
$u_{0}^{(\mathrm{i})}=\frac{1}{A}\left[1-R\mathcal{K}_{1}\left(R\right)\mathcal{I}_{0}\left(r\right)\right],$ (91)
$u_{0}^{(\mathrm{o})}=\frac{1}{A}\mathcal{I}_{1}\left(R\right)\mathcal{K}_{0}\left(r\right),$ (92)
$u_{1}^{(\mathrm{i})}=\frac{R}{2A}\left[r\mathcal{K}_{1}\left(R\right)\mathcal{I}_{0}\left(r\right)-R\mathcal{K}_{2}\left(R\right)\mathcal{I}_{1}\left(r\right)\right]\cos\theta,$ (93)
$u_{1}^{(\mathrm{o})}=\frac{R}{2A}\left[-r\mathcal{I}_{1}\left(R\right)\mathcal{K}_{0}\left(r\right)+R\mathcal{I}_{2}\left(R\right)\mathcal{K}_{1}\left(r\right)\right]\cos\theta,$ (94)
$u_{2}^{(\mathrm{i})}=$ $\frac{R^{2}}{32A}\left[r^{2}\left(\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)-\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right)-2\left(R\mathcal{K}_{1}(R)\mathcal{I}_{0}(r)-r\mathcal{K}_{0}(R)\mathcal{I}_{1}(r)\right),\right]$
$+\frac{R^{2}}{64A}\left[2r^{2}\left(\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)-\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right)-Rr\left(\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)\right)\right]\cos 2\theta,$ (95)
$u_{2}^{(\mathrm{o})}=$ $\frac{R^{2}}{32A}\left[r^{2}\left(\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)-\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right)+2\left(R\mathcal{I}_{1}(R)\mathcal{K}_{0}(r)-r\mathcal{I}_{0}(R)\mathcal{K}_{1}(r)\right)\right]$
$+\frac{R^{2}}{64A}\left[2r^{2}\left(\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)-\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right)-Rr\left(\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)\right)\right]\cos 2\theta,$ (96)
$u_{3}^{(\mathrm{i})}=$ $\frac{R^{2}}{128A}\left[r(R^{2}+r^{2})\left(\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)-\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)\right)+4r\left(R\mathcal{K}_{1}(R)\mathcal{I}_{0}(r)-r\mathcal{K}_{0}(R)\mathcal{I}_{1}(r)\right)\right]\cos\theta$
$+\frac{R^{2}}{1152A}\left[-3r^{3}\left(\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)-\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right)+3r^{2}R\left(\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)\right)\right$.
$\left.\qquad-R^{2}r\left(\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)-\mathcal{K}_{4}(R)\mathcal{I}_{4}(r)\right)\right]\cos 3\theta,$ (97)
$u_{3}^{(\mathrm{o})}=$ $\frac{R^{2}}{128A}\left[r(R^{2}+r^{2})\left(\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)-\mathcal{K}_{0}(R)\mathcal{K}_{0}(r)\right)-4r\left(R\mathcal{I}_{1}(R)\mathcal{K}_{0}(r)-r\mathcal{I}_{0}(R)\mathcal{K}_{1}(r)\right)\right]\cos\theta$
$+\frac{R^{2}}{1152A}\left[-3r^{3}\left(\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)-\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right)+3r^{2}R\left(\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)\right)\right$.
$\left.\qquad-R^{2}r\left(\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)-\mathcal{I}_{4}(R)\mathcal{K}_{4}(r)\right)\right]\cos 3\theta,$ (98)
$\tilde{u}^{(i)}_{0}$= $\frac{1}{A}\left[R^{2}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right]\left(\tilde{a}_{2}\cos 2\theta+\tilde{b}_{2}\sin 2\theta\right),$ (99)
$\tilde{u}^{(\mathrm{o})}_{0}=\frac{1}{A}\left[R^{2}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right]\left(\tilde{a}_{2}\cos 2\theta+\tilde{b}_{2}\sin 2\theta\right),$ (100)
$\tilde{u}^{(\mathrm{i})}_{1}$= $\frac{R^{2}}{4A}\left[-r\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)+R\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)\right]\left(\tilde{a}_{2}\cos\theta+\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{4A}\left[-r\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)+R\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)\right]\left(\tilde{a}_{2}\cos 3\theta+\tilde{b}_{2}\sin 3\theta\right),$ (101)
$\tilde{u}^{(\mathrm{o})}_{1}$= $\frac{R^{2}}{4A}\left[-r\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)+R\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)\right]\left(\tilde{a}_{2}\cos\theta+\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{4A}\left[-r\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)+R\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)\right]\left(\tilde{a}_{2}\cos 3\theta+\tilde{b}_{2}\sin 3\theta\right),$ (102)
$\tilde{u}^{(\mathrm{i})}_{2}$= $\frac{R^{2}}{32A}\left[R^{2}\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)-2rR\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)+r^{2}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right]\tilde{a}_{2}$
$+\frac{R^{2}}{16A}\left[-\frac{3}{2}rR\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)+(r^{2}+R^{2})\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)-\frac{1}{2}rR\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)\right]\left(\tilde{a}_{2}\cos 2\theta+\tilde{b}_{2}\sin 2\theta\right)$
$+\frac{R^{2}}{32A}\left[r^{2}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)-2rR\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)+R^{2}\mathcal{K}_{4}(R)\mathcal{I}_{4}(r)\right]\left(\tilde{a}_{2}\cos 4\theta+\tilde{b}_{2}\sin 4\theta\right),$ (103)
$\tilde{u}^{(\mathrm{o})}_{2}$= $\frac{R^{2}}{32A}\left[R^{2}\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)-2rR\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)+r^{2}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right]$
$+\frac{R^{2}}{16A}\left[-\frac{3}{2}rR\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)+(r^{2}+R^{2})\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)-\frac{1}{2}rR\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)\right]\left(\tilde{a}_{2}\cos 2\theta+\tilde{b}_{2}\sin 2\theta\right)$
$+\frac{R^{2}}{32A}\left[r^{2}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)-2rR\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)+R^{2}\mathcal{I}_{4}(R)\mathcal{K}_{4}(r)\right]\left(\tilde{a}_{2}\cos 4\theta+\tilde{b}_{2}\sin 4\theta\right),$ (104)
$\tilde{u}^{(\mathrm{i})}_{3}$= $\frac{R^{2}}{384A}\left[R^{3}\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-3rR^{2}\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)+3Rr^{2}\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-r^{3}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right]\left(\tilde{a}_{2}\cos\theta-\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{128A}\left[-3rR^{2}\mathcal{K}_{0}(R)\mathcal{I}_{0}(r)+R(3r^{2}+R^{2})\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-r^{3}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)\right]\left(\tilde{a}_{2}\cos\theta+\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{128A}\left[2r^{2}R\mathcal{K}_{1}(R)\mathcal{I}_{1}(r)-r\left(r^{2}+\frac{8}{3}R^{2}\right)\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)+R(r^{2}+R^{2})\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)-\frac{1}{3}rR^{2}\mathcal{K}_{4}(R)\mathcal{I}_{4}(r)\right]$
$\quad\times\left(\tilde{a}_{2}\cos 3\theta+\tilde{b}_{2}\sin 3\theta\right)$
$+\frac{R^{2}}{384A}\left[-r^{3}\mathcal{K}_{2}(R)\mathcal{I}_{2}(r)+3r^{2}R\mathcal{K}_{3}(R)\mathcal{I}_{3}(r)-3rR^{2}\mathcal{K}_{4}(R)\mathcal{I}_{4}(r)+R^{3}\mathcal{K}_{5}(R)\mathcal{I}_{5}(r)\right]\left(\tilde{a}_{2}\cos 5\theta+\tilde{b}_{2}\sin 5\theta\right),$ (105)
$\tilde{u}^{(\mathrm{o})}_{3}$= $\frac{R^{2}}{384A}\left[R^{3}\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-3rR^{2}\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)+3Rr^{2}\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-r^{3}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right]\left(\tilde{a}_{2}\cos\theta-\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{128A}\left[-3rR^{2}\mathcal{I}_{0}(R)\mathcal{K}_{0}(r)+R(3r^{2}+R^{2})\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-r^{3}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)\right]\left(\tilde{a}_{2}\cos\theta+\tilde{b}_{2}\sin\theta\right)$
$+\frac{R^{2}}{128A}\left[2r^{2}R\mathcal{I}_{1}(R)\mathcal{K}_{1}(r)-r\left(r^{2}+\frac{8}{3}R^{2}\right)\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)+R(r^{2}+R^{2})\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)-\frac{1}{3}rR^{2}\mathcal{I}_{4}(R)\mathcal{K}_{4}(r)\right]$
$\quad\times\left(\tilde{a}_{2}\cos 3\theta+\tilde{b}_{2}\sin 3\theta\right)$
$+\frac{R^{2}}{384A}\left[-r^{3}\mathcal{I}_{2}(R)\mathcal{K}_{2}(r)+3r^{2}R\mathcal{I}_{3}(R)\mathcal{K}_{3}(r)-3rR^{2}\mathcal{I}_{4}(R)\mathcal{K}_{4}(r)+R^{3}\mathcal{I}_{5}(R)\mathcal{K}_{5}(r)\right]\left(\tilde{a}_{2}\cos 5\theta+\tilde{b}_{2}\sin 5\theta\right).$ (106)
Appendix C Fitting of the coefficient of the cubic term of the deformation
Since our model started from the free energy considering the first order of the deformation magnitude, we cannot analytically derive the coefficient of the cubic term with respect to the deformation magnitude. For the ODE approach, we need to know the value of the coefficient of the cubic term. Therefore, we ran the numerical simulation and obtained the coefficient by fitting the simulation data. Since we only needed to consider the transition from an IC droplet to an ID droplet, we did not include the dynamics of the translational motion but only considered the deformation dynamics coupled with the time evolution of the concentration field. Other parameters were set to be the same as described in the main text. The plot of the deformation magnitude $s$ versus $\kappa_{2}$ is shown with red dots in Fig. 5.
Figure 5: Magnitude $s$ of deformation depending on $\kappa_{2}$ obtained by the numerical simulation only considering the deformation dynamics. The simulation results are shown with red dots. The result of the fitting to a function $\kappa_{2}=p_{0}-p_{2}s^{2}$ is shown with a blue line. Here, $p_{0}$ and $p_{2}$ were obtained to be $p_{0}=0.1184$ and $p_{2}=0.1436$ from the fitting. The green broken line shows the curve using the analytically estimated value for $\tilde{g}_{0}$.
The simulation results are fitted by the function
$\kappa_{2}=p_{0}-p_{2}s^{2},$ (107)
with the least square method. Since we needed to focus on the region with small $\left|s\right|$, we only used the data with $0.001<\left|s\right|<0.3$. By the fitting, we obtained $p_{0}=0.1184$ and $p_{2}=0.1436$. The resulting curve is shown with a blue solid line in Fig. 5. From the theoretical analysis, the constant term corresponding to $p_{0}$ is estimated to be $\tilde{g}_{0}$, which is estimated as $\tilde{g}_{0}\simeq 0.1196$ with the corresponding parameters. The curve using this estimated value for $\tilde{g}_{0}$ is also plotted by a green broken line in Fig. 5. The difference between the estimated value of $\tilde{g}_{0}$ and $p_{0}$ may be due to the smoothing effect. For obtaining the phase diagram in Fig. 4, we adopted the estimated value from the theory for $\tilde{g}_{0}$ and $p_{2}\simeq 0.1436$ for $h_{0}$.
Appendix D Details on analyses of the ODE model
First, we discuss the stability of an immobile circular (IC) droplet, $v=s=0$. Considering the linear terms of Eqs. (67) and (69), the IC droplet is stable when $\eta_{t}>f_{1}$ and $\kappa_{2}>\tilde{g}_{0}$. In the case of $\kappa_{2}<\tilde{g}_{0}$, the solution corresponding to an immobile deformed (ID) droplet exists: $v=0$ and $s=\sqrt{(\tilde{g}_{0}-\kappa_{2})/h_{0}}$. This transition from IC to ID droplets is a supercritical pitchfork bifurcation.
Next, we consider a moving deformed (MD) droplet, in which both $v$ and $s$ have finite values. From Eqs. (68) and (70), we obtain the equation for $\xi=\psi-\phi$ as
$\frac{d\xi}{dt}=\left(\frac{g_{2}v^{2}}{2\eta_{2}s}+\frac{\tilde{f}_{1}s}{m}\right)\sin 2\xi.$ (108)
Considering that the coefficient of $\sin 2\xi$ is always positive, $\xi$ converges to $\pm\pi/2$. Thus, hereafter, we assume $\xi=\pm\pi/2$ always hold. It is notable that $\xi=\pm\pi/2$ indicates that the droplet moves its minor-axis direction. Then, the solution corresponding to an MD droplet is obtained by simultaneously solving
$\left(-\eta_{t}+f_{1}\right)v+\tilde{f}_{1}sv-f_{3}v^{3}=0,$ (109)
$\left(-\kappa_{2}+\tilde{g}_{0}\right)s+g_{2}v^{2}-\tilde{g}_{2}sv^{2}-h_{0}s^{3}=0.$ (110)
By eliminating $v$ from these two equations, we obtain the equation for $s$
$H(s)=$ $f_{3}h_{0}s^{3}+\tilde{f}_{1}\tilde{g}_{2}s^{2}+\left[-f_{3}(\tilde{g}_{0}-\kappa_{2})\right$.
$\left.-\tilde{g}_{2}(\eta_{t}-f_{1})-g_{2}\tilde{f}_{1}\right]s+g_{2}(\eta_{t}-f_{1})=0.$ (111)
The number of solutions of Eq. (111) changes when $H(s)$ touches the $s$ axis. Thus, by obtaining $s$ that holds
$H(s)=\frac{dH}{ds}=0,$ (112)
we can obtain the boundary of the region in which the solution for an MD droplet exists.
Finally, the stability of an ID droplet is discussed. By substituting $s$ in Eq. (67) with the value of $s$ for an ID droplet, we obtain
$m\frac{dv}{dt}=\left(-\eta_{t}+f_{1}+\tilde{f}_{1}\sqrt{\frac{\tilde{g}_{0}-\kappa_{2}}{h_{0}}}\right)v-f_{3}v^{3}.$ (113)
Therefore, the threshold of the stability for an ID droplet is
$\eta_{t}=f_{1}+\tilde{f}_{1}\sqrt{\frac{\tilde{g}_{0}-\kappa_{2}}{h_{0}}}.$ (114)
Therefore, the curves $\eta_{t}=f_{1}$, $\kappa_{2}=\tilde{g}_{0}$, Eq. (112) with Eq. (111), and Eq. (114) indicate the boundary of a stable IC droplet region when decreasing $\eta_{t}$, the boundary between stable IC and ID droplet regions, the boundary of a stable MD droplet when increasing $\eta_{t}$ or $\kappa_{2}$, and the boundary of the stable ID droplet when decreasing $\eta_{t}$ or $\kappa_{2}$, respectively, as shown in Fig. 4.